Evaluate the ∫y(y²-2)²dy


1
Expert's answer
2021-02-24T12:13:52-0500

Solution

∫y(y²-2)²dy = (1/2) ∫(y²-2)²dy2

Substitution x = y2 => ∫y(y²-2)²dy = (1/2) ∫(x-2)²dx = (1/2) ∫(x-2)² d(x-2) = (x-2)3 /6 + C =   (y2-2)3 /6 + C

Answer

∫y(y²-2)²dy = (y2-2)3 /6 + C